Please help me. I am having problem understanding this question.

: If your test score are 80 and 100, then the arithmetic mean of your scores is 90. The geometric mean of the scores is a number x such that
80/x = x/100
Are you better off with the arithmetic mean or the geometric mean?
Thank you for your help.

Answer by stanbon(75887)

If your test score are 80 and 100, then the arithmetic mean of your scores is 90. The geometric mean of the scores is a number x such that
80/x = x/100
Are you better off with the arithmetic mean or the geometric mean?
---------------
Find the geometric mean:
80/x = x/100
x^2 = 80*100
x = 89.443
==============
Cheers,
Stan H.